What are the steps to solving the system
x^2+y^2=169
x-y=7

Answer by jojo14344(1513)

Okay, x^2+y^2=169, mark this equation 1.
Also, x-y=7, mark this then equation 2.
In equation 2 we get, y=x-7, right? Substitute this y value in equation 1,
Therefore,
x^2+(x-7)^2=169
x^2+x^2-14x+49=169
2x^2-14x-120=0
(2x+10)(x-12)=0
It has 2 values, first
x=-10/2=-5,
For y, go back equation 2, so y =-5-7=-12
This equation has (-5,-12)
Also the other value x=12, and y=12-7=5
For this equation has (12,5)
Thank you,
jojo